OpenStudy (anonymous):

Can some one please help me with 4 questions? Will fan and medal :)

Nnesha (nnesha):

O_O question

OpenStudy (anonymous):

sorry my computer was lagging :p Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

none of these "show" anything, but \(1+1=2\) is an EXAMPLE of them being closed

OpenStudy (anonymous):

sorry my laptop is lagging really bad! And all i see is question marks :/

OpenStudy (anonymous):

Choose all answers that are correct. 3. (1 pt) Which sets of numbers are closed under division? Choose all answers that are correct. A. rational numbers B. integers C. {–1, 0, 1} D. whole numbers